# rowsep=0pt

### v1

@@ 2 Related work
X-vector–based speaker anonimization - The original x-vector–based pipeline introduced in [3] is the basis of much of the related work reported recently. An approach to dispense with the intermediate acoustic model was proposed in [2]. More refined x-vector anonymization functions were proposed in later work [10, 11], with some [12] achieving notable improvements to privacy protection levels, albeit under the assumption that the attacker does not have full knowledge of the anonymization system. Whatever the approach, x-vector perturbation does not prevent speaker-related information contained in the F0 curve and linguistic features from leaking into the anonymized waveform upon vocoding. As a result, the x-vector which can be re-extracted from the anonymized waveform by a privacy adversary who wishes to reidentify the speaker tends to drift away from that at the vocoder input [13]. While the drift can be beneficial to anonymization, it hinders the design of more effective anonymization functions and can also be inverted by an adversary to undo the privacy protection [14]. Attempts to sanitize speaker information from linguistic features have also been explored, e.g. [7] based on the concept of differential privacy, which reports improvements to privacy at the cost of reduced utility and pitch correlation. The same issue was tackled in [5] by means of feature quantization, which is shown to be effective as a speaker information bottleneck, though with a degradation to utility. In this paper, we follow a similar approach, but propose a completely new synthesis pipeline in which we avoid the leakage of information between different speech components by design.
@@
NAC language modeling - NACs were proposed recently for audio compression [15, 16]. They consist in convolutional autoencoders that compress audio to low-bitrate, tokenized representations which support high-fidelity decoding. Due to their discretized nature, encoded representations can be modeled using techniques normally used for language-related tasks, such as transformers. This idea was first introduced in [8] for a variety of audio generation tasks, and tailored to text-to-speech (TTS) in [9]. A transformer is used to convert the input (be it audio or text) to a set of high-level semantic tokens. These are then fed to another transformer which converts them into NAC acoustic tokens which can be decoded to resynthesize an audio signal.
@@
The same technique can also be applied to voice conversion [8, 17], and is hence ideally suited to speaker anonymization. NAC language models operate on quantized codes, which are known to be beneficial to privacy protection [5]. Moreover, such models appear to naturally disentangle linguistic information into semantic tokens, while encoding speaker information and recording conditions mostly in acoustic tokens [8]. Hence, we propose a speaker anonymization system whereby an input utterance is re-synthesized by means of a NAC language model. Semantic tokens are kept unchanged, while acoustic tokens are substituted with those of a different speaker, the goal being to preserve the linguistic content of the signal while suppressing information related to the original speaker.
@@ 3.2 Anonymization technique
A set of semantic tokens $\mathbf{s}$ is first extracted from the input utterance. They encode the high-level spoken content. Their quantization helps to suppress speaker-related information.
A pseudo-speaker is chosen by randomly selecting an acoustic prompt $\mathbf{\tilde{\mathbf{a}}}$ from the speaker prompt pool. Anonymization can be performed at either speaker or utterance levels. At the speaker level, anonymization is performed using the same speaker prompt for each utterance corresponding to any one speaker. In contrast, for utterance level anonymization, a speaker prompt is selected at random for each utterance. While several anonymization systems include techniques to synthesize fictitious voices [5, 10, 11, 12], here we use real voices as pseudo-speakers to focus our analysis on the intrinsic anonymization capability of the NAC language model.
@@
EER (%) WER (%) $G_{VD}$ $\boldsymbol{\rho}^{F_{0}}$ EER (%) WER (%) $G_{VD}$ $\boldsymbol{\rho}^{F_{0}}$
@@
Original (eval. pipeline of [18]) 1.5 2.5 n.a. 1.1 7.6 n.a.
@@ 4 Experimental setup
Our codebase is branched from Bark, [^2] The original source code is available at www.github.com/suno-ai/bark, though we built our system from the port included in the CoquiTTS library available at www.github.com/coqui-ai/TTS. Our source code, as well as audio samples, will be made available upon publication. an open-source, NAC-based TTS system which is very similar to VALL-E [9]. The modules described in Section 3.1 are all taken from Bark. The semantic encoder has a HuBERT backbone [19] and a LSTM [20] back-end which predicts the semantic token associated to the HuBERT feature vector output at each frame. The semantic dictionary is of size $N_{S}=10000$. The coarse and fine transformers are 12-layer GPT-like models [21], each having $Q=8$ different codebooks with $N_{Q}=1024$ codewords. The first $Q_{C}=2$ codebooks are considered coarse. The NAC is EnCodec [16]. The difference between Bark and our system is that, being a TTS model, Bark estimates semantic tokens $\mathbf{s}$ corresponding to an input text using a further semantic transformer. In our implementation, we bypass the semantic transformer and use ground truth semantic tokens from the input waveform thereby providing for voice conversion instead of TTS. With this setup, we are able to use pre-trained Bark modules off-the-shelf, without the need for any training.
@@
We adopt the Voice Privacy Challenge 2022 protocol [2] for evaluation. The test set comprises subsets of the LibriSpeech [22] and VCTK [23] databases. The pool of speaker prompts is taken from the Bark voice library. It consists of 130 utterances collected from speakers of different gender and nationality.333www.github.com/suno-ai/bark/tree/main/bark/assets/prompts/v2 The threat model is the semi-informed attack described in [2]. Trial utterances are anonymized at the speaker level. The attacker is assumed to have access to the anonymization system. They anonymize a set of external data (librispeech-clean-360) at the utterance level and use it to train an ASV system (a TDNN with a PLDA back-end [24]). They also have access to original (non-protected) enrollment utterances which they anonymize at the speaker level. The attacker thus has enrollment and trial utterances both of which are anoymized and uses an ASV system to verify whether they correspond to the same speaker. The attacker has no knowledge of which pseudo-speaker was used for anonymization on the test utterance and will hence likely select a different pseudo-speaker to anonymize the enrollment utterance. The privacy metric is the resulting EER estimated from a large number of ASV trials. Utility is assessed by training an ASR system on the same anonymized version of librispeech-clean-360 and by estimating the word error rate (WER) from anonymized test data. Two additional metrics are defined in the VoicePrivacy Challenge evaluation plan [2]. The first is the F0 curve correlation $\boldsymbol{\rho}^{F0}$ between original and anonymized utterance which is used as a measure of prosody preservation. The second is the gain of voice distinctiveness $G_{VD}$ which is used to estimate how well the anonymized voices of different speakers can be distinguished [2].
@@ 5 Results
In an attempt to shed light on the cause for this phenomenon, we repeated similar experiments using a different ASR architecture, namely that reported in [18], which is retrained according to the same setup described in Section 4. The issue persists. The WER increases from 2.5% to 4.6% for the LibriSpeech subset and from 7.6% to 15.5% for the VCTK subset. These findings suggest that the degradation to utility is more dependend on the NAC language model than on the ASR system. As suggested in [8], this could be due to the quality of some pseudo-speaker prompts, since the extracted fine acoustic tokens tend also to capture aspects of the (potentially poor) recording conditions, the characteristics of which are then transferred to anonymized outputs. More thorough experimentation to help us better understand this phenomena is already underway.

### v3

@@ 2 Related work
X-vector–based speaker anonymization - The original x-vector–based pipeline introduced in [3] is the basis of much of the related work reported recently. An approach to dispense with the intermediate acoustic model was proposed in [2]. More refined x-vector anonymization functions were proposed in later work [10, 11], with some [12] achieving notable improvements to privacy protection levels, albeit under the assumption that the attacker does not have full knowledge of the anonymization system. Whatever the approach, x-vector perturbation does not prevent speaker-related information contained in the F0 curve and linguistic features from leaking into the anonymized waveform upon vocoding. As a result, the x-vector which can be re-extracted from the anonymized waveform by a privacy adversary who wishes to re-identify the speaker tends to drift away from that at the vocoder input [13]. While the drift can be beneficial to anonymization, it hinders the design of more effective anonymization functions and can also be inverted by an adversary to undo the privacy protection [14]. Attempts to sanitize speaker information from linguistic features have also been explored, e.g. [7] based on the concept of differential privacy, which reports improvements to privacy at the cost of reduced utility and pitch correlation. The same issue was tackled in [5] by means of feature quantization, which is shown to be effective as a speaker information bottleneck, though with a degradation to utility. In this paper, we follow a similar approach, but propose a completely new synthesis pipeline in which we avoid the leakage of information between different speech components by design.
@@
NAC language modeling - NACs were proposed recently for audio compression [15, 16]. They consist in convolutional autoencoders that compress audio to low-bitrate, tokenized representations which support high-fidelity decoding. Due to their discretized nature, encoded representations can be modeled using techniques normally used for language-related tasks, such as transformers. This idea was first introduced in [8] for a variety of audio generation tasks, and tailored to text-to-speech (TTS) in [9]. A transformer is used to convert the input (be it audio or text) to a set of high-level semantic tokens. These are then fed to another transformer which converts them into NAC acoustic tokens which can be decoded to re-synthesize an audio signal.
@@
The same technique can also be applied to voice conversion [8, 17], and is hence ideally suited to speaker anonymization. NAC language models operate on quantized codes, which are known to be beneficial to privacy protection [5]. Moreover, such models appear to naturally disentangle linguistic information into semantic tokens, while encoding speaker information and recording conditions mostly in acoustic tokens [8]. Hence, we propose a speaker anonymization system whereby an input utterance is re-synthesized by means of a NAC language model. Semantic tokens are kept unchanged, while acoustic tokens are substituted with those of a different speaker: the goal is to preserve the linguistic content of the signal while suppressing information related to the original speaker.
@@ 3.2 Anonymization technique
A set of semantic tokens $\mathbf{s}$ is first extracted from the input utterance. These tokens encode the high-level spoken content. Their quantization helps to suppress speaker-related information.
A pseudo-speaker is chosen by randomly selecting an acoustic prompt $\mathbf{\tilde{\mathbf{a}}}$ from the speaker prompt pool. Anonymization can be performed at either speaker or utterance levels. At the speaker level, anonymization is performed using the same speaker prompt for each utterance corresponding to one specific speaker. In contrast, for utterance level anonymization, a speaker prompt is selected at random for each utterance. While several anonymization systems include techniques to synthesize fictitious voices [5, 10, 11, 12], here we use real voices as pseudo-speakers to focus our analysis on the intrinsic anonymization capability of the NAC language model.
@@
EER (%) $\uparrow$ WER (%) $\downarrow$ $G_{VD}\uparrow$ $\boldsymbol{\rho}^{F_{0}}\uparrow$ EER (%) $\uparrow$ WER (%) $\downarrow$ $G_{VD}\uparrow$ $\boldsymbol{\rho}^{F_{0}}\uparrow$
@@
Original (eval. pipeline of [18]) 1.5 2.5 1.1 7.6
@@ 4 Experimental setup
Our codebase is branched from Bark, [^2] The original source code is available at www.github.com/suno-ai/bark, though we built our system from the port included in the CoquiTTS library available at www.github.com/coqui-ai/TTS. Our source code, as well as audio samples, are available at www.github.com/eurecom-asp/spk_anon_nac_lm. an open-source, NAC-based TTS system which is very similar to VALL-E [9]. The modules described in Section 3.1 are all taken from Bark. The semantic encoder has a HuBERT backbone [19] and a LSTM [20] back-end which predicts the semantic token associated to the HuBERT feature vector output at each frame. The semantic dictionary is of size $N_{S}=10000$. The coarse and fine transformers are 12-layer GPT-like models [21], each having $Q=8$ different codebooks with $N_{Q}=1024$ codewords. The first $Q_{C}=2$ codebooks are considered coarse. The NAC is EnCodec [16]. The difference between Bark and our system is that, being a TTS model, Bark estimates semantic tokens $\mathbf{s}$ corresponding to an input text using a further semantic transformer. In our implementation, we bypass the semantic transformer and use ground truth semantic tokens from the input waveform thereby providing for voice conversion instead of TTS. With this setup, we are able to use pre-trained Bark modules off-the-shelf, without the need for any training.
@@
We adopt the Voice Privacy Challenge 2022 protocol [2] for evaluation. The test set comprises subsets of the LibriSpeech [22] and VCTK [23] databases. The pool of speaker prompts is taken from the Bark voice library. It consists of 130 utterances collected from speakers of different gender and nationality.333www.github.com/suno-ai/bark/tree/main/bark/assets/prompts/v2 The threat model is the semi-informed attack described in [2]. Trial utterances are anonymized at the speaker level. The attacker is assumed to have access to the anonymization system. They anonymize a set of external data (librispeech-clean-360) at the utterance level and use it to train an ASV system (a TDNN with a PLDA back-end [24]). They also have access to original (non-protected) enrollment utterances which they anonymize at the speaker level. The attacker thus has enrollment and trial utterances both of which are anonymized and uses an ASV system to verify whether they correspond to the same speaker. The attacker has no knowledge of which pseudo-speaker was used for anonymization on the test utterance and will hence likely select a different pseudo-speaker to anonymize the enrollment utterance. The privacy metric is the resulting EER estimated from a large number of ASV trials; a higher EER corresponds to better privacy. Utility is assessed by training an ASR system on the same anonymized version of librispeech-clean-360 and by estimating the word error rate (WER) from anonymized test data; a lower WER corresponds to better utility. Two additional metrics are defined in the VoicePrivacy Challenge evaluation plan [2]. The first is the F0 curve correlation $\boldsymbol{\rho}^{F0}$ between original and anonymized utterance which is used as a measure of prosody preservation. The second is the gain of voice distinctiveness $G_{VD}$ which is used to estimate how well the anonymized voices of different speakers can be distinguished [2].
@@ 5 Results
In an attempt to shed light on the cause for this phenomenon, we repeated similar experiments using a different ASR architecture, namely that reported in [18], which is retrained according to the same setup described in Section 4. The issue persists. The WER increases from 2.5% to 4.6% for the LibriSpeech subset and from 7.6% to 15.5% for the VCTK subset. These findings suggest that the degradation to utility is more dependent on the NAC language model than on the ASR system. As suggested in [8], this could be due to the quality of some pseudo-speaker prompts, since the extracted fine acoustic tokens tend also to capture aspects of the (potentially poor) recording conditions, the characteristics of which are then transferred to anonymized outputs. More thorough experimentation to help us better understand this phenomena is already underway.